# LLMs Reproduce Stereotypes of Sexual and Gender Minorities

### v2

@@ 5 SCM Axes in LLM Text Generation
Partly inspired by Cheng et al. (2023), we used five prompts (Table 4) to simulate a creative text generation task focused on a member of a specific group. The prompts are designed so that the results highlight both general personality characteristics of individuals from these groups and these characteristics in a specific, action-oriented setting, namely a narrative. We ran all prompts 5 times in total, giving 50 outputs per group.
@@
Table 4: Our text generation prompts. The first three follow Cheng et al. (2023).
@@ 5.1 Results
Text generation is more open-ended than even the elicitation task, so we need a way to focus on descriptive words. We used SpaCy Honnibal and Montani (2017) to identify nouns, adjectives, and verbs in generated texts, focusing solely on these words in subsequent analyses. Quantitative results (Figure 5) are not as consistent with the human survey as earlier results, but we still observe that the relative associations of Competence track those of the survey. Results are also consistent with the survey in the sense that they prefer Warmth to Competence for most groups.
001010101020202020GPTof termsCompetenceWarmth
001010101020202020Geminiof terms
womenmennonbinarygaylesbianbisexualheterosexual001010101020202020LLaMAof terms
Figure 5: Percentage of Warmth and Competence terms obtained by LLM text generation, using the method of Section 4.2 to associate words to concepts. Since unrelated terms occur, percentages do not sum to 100.
@@
GPT individual presence connection love mountains family dragon
@@
Gemini gender woman music strength woman power knowledge
@@
LLaMA gender love sexuality love equality self man
@@
To focus on words that the models strongly associate with each group, we borrow an idea from Wan et al. (2023), and compute the Odds Ratio (OR) for each word and group. OR is the ratio of two conditional probabilities: that of generating a word conditioned on the group of interest, and that of generating the same word, not conditioned on any group. The results (Table 5) show that generated texts reinforce the Warmth-Competence stereotypes found throughout our results. For example, stories about women focus on Warmth (e.g. “ability to heal others", “a passionate advocate for social justice"); and about men on Competence (e.g. a man learning to rock climb grows “stronger and more confident"). A nonbinary person “often felt misunderstood" by others and “whispers and sideways glances" followed them. A lesbian faces “discrimination and marginalization" throughout their life. A bisexual is “condemned" and called “a deviant, a threat to the village’s morals." These passages emphasize marginalization and pain for minoritized identities, reifying painful experience as most representative of their lives even in creative stories. Indeed, all LLMs frequently generated words about struggle (e.g., “challenges”, “justice”, “messy”) for nonbinary people, bisexuals, and lesbians, a pattern also observed by Dhingra et al. (2023).
@@ 6 Conclusion
Our survey does not ask participants about their personal opinion. This is the established approach to SCM survey methods Fiske et al. (2002); Cuddy et al. (2007), and makes our results more amenable to comparisons with that work. However, it is known that people’s perceptions may be biased by exposure to news coverage. For example, after a rare disaster occurs (such as a plane crash) survey respondents are likely to over-estimate the risk of such a disaster. Similarly, it’s possible that in the political climate of summer 2024, when we conducted our survey, respondents have a biased perspective of how society views sexual and gender minorities, based on widespread and polarized political news coverage in several English-speaking countries. In contrast, the training data of the LLMs we study likely almagamates attitudes across many years, rather than at a specific point in time.
@@ Acknowledgments
We thank Fengyu Liu and Yuanqi Shi for helpful discussion of this work; and Sharon Goldwater, Coleman Haley, Oli Liu, Yen Meng, Sung-Lin Yeh, and anonymous TACL reviewers for comments on earlier drafts of this paper.

### v3

@@ 5 SCM Axes in LLM Text Generation
Partly inspired by Cheng et al. (2023), we used five prompts (Table˜4) to simulate a creative text generation task focused on a member of a specific group. The prompts are designed so that the results highlight both general personality characteristics of individuals from these groups and these characteristics in a specific, action-oriented setting, namely a narrative. We ran all prompts 5 times in total, giving 50 outputs per group.
@@
Table 4: LLM text generation prompts. The first three follow Cheng et al. (2023).
@@ 5.1 Results
Text generation is more open-ended than even the elicitation task, so we need a way to focus on descriptive words. We used SpaCy Honnibal and Montani (2017) to identify nouns, adjectives, and verbs in generated texts, focusing solely on these words in subsequent analyses. Quantitative results (Figure˜5) are not as consistent with the human survey as earlier results, but we still observe that the relative associations of Competence track those of the survey. Results are also consistent with the survey in the sense that they prefer Warmth to Competence for most groups.
010102020GPT% of termsCompetenceWarmth
010102020Gemini% of terms
womenmennonbinarygaylesbianbisexualheterosexual010102020LLaMA% of terms
Figure 5: Percentage of Warmth and Competence terms obtained by LLM text generation, using the method of Section˜4.2 to associate words to concepts. Since unrelated terms occur, percentages do not sum to 100.
@@
GPT
individual presence connection love mountains family dragon
@@
Gemini
gender woman music strength woman power knowledge
@@
LLaMA
gender love sexuality love equality self man
@@
To focus on words that the models strongly associate with each group, we borrow an idea from Wan et al. (2023), and compute the Odds Ratio (OR) for each word and group. OR is the ratio of two conditional probabilities: that of generating a word conditioned on the group of interest, and that of generating the same word, not conditioned on any group. The results (Table˜5) show that generated texts reinforce the Warmth-Competence stereotypes found throughout our results. For example, stories about women focus on Warmth (e.g. “ability to heal others", “a passionate advocate for social justice"); and about men on Competence (e.g. a man learning to rock climb grows “stronger and more confident"). A nonbinary person “often felt misunderstood" by others and “whispers and sideways glances" followed them. A lesbian faces “discrimination and marginalization" throughout their life. A bisexual is “condemned" and called “a deviant, a threat to the village’s morals." These passages emphasize marginalization and pain for minoritized identities, reifying painful experience as most representative of their lives even in creative stories. Indeed, all LLMs frequently generated words about struggle (e.g., “challenges”, “justice”, “messy”) for nonbinary people, bisexuals, and lesbians, a pattern also observed by Dhingra et al. (2023), and which Ungless et al. (2025) include in a community-centered taxonomy of harms that LLMs pose towards nonbinary individuals.
@@ 6 Conclusion
Our survey inherits limitations of the SCM framework. For example, the SCM theory centers Warmth and Competence as attributes that capture stereotypical associations across many contexts; this makes it widely applicable, but also makes it reductive, ignoring features of stereotypes that may be of interest in specific contexts such as ours, perhaps only in those contexts. This reflects an unavoidable tension in socially oriented research, between the general and the specific. While we chose the SCM because it has been widely validated, other choices could be explored.
A second limitation that we inherit from the SCM is that we do not ask participants about their personal opinion, following established SCM survey methods Fiske et al. (2002); Cuddy et al. (2007). This makes our results more amenable to comparisons with that work, but it is known that people’s perceptions may be biased by exposure to news coverage. For example, after a rare disaster occurs (such as a plane crash) survey respondents are likely to over-estimate the risk of such a disaster. Similarly, it’s possible that in the political climate of summer 2024, when we conducted our survey, respondents have a biased perspective of how society views sexual and gender minorities, based on widespread and polarized political news coverage in several English-speaking countries. In contrast, the training data of the LLMs we study likely almagamates attitudes across many years, rather than at a specific point in time. It has been previously shown that neural network language models (of which LLMs are one type) can capture stereotypes that reflect the time period of the corpora they are trained on (Garg et al., 2018).
@@ Acknowledgments
We thank Fengyu Liu and Yuanqi Shi for helpful discussion of this work; and Sharon Goldwater, Coleman Haley, Oli Liu, Yen Meng, Sung-Lin Yeh, and anonymous TACL and ARR reviewers and area chairs for constructive comments of earlier drafts of this paper.
Appendix A Additional Prompts
Rating
I am studying how different groups are perceived. I am not interested in your personal beliefs, but in how you think they are viewed by others. Please make a rating using a 5-point scale (1=not at all to 5=extremely) of how members of [group] are viewed by society. Again, I am not interested in your personal beliefs, but in how you think society views them.
Selection
Choose an attribute from the following list for a person from the [group] sexual orientation/gender based on societal views. Choose only one. Provide a rating using a 5-point scale (1=not at all to 5=extremely), on the basis of how much the group is viewed this way by society. [list of all Warmth-Competence terms]
Table 6: Additional LLM prompts to those given in Table 2, showing variables in teal that are sourced from the lists in Table 1.
Appendix B Analysis of Refusals
For SCM prompts, we define refusal as a failure to output a score or keyword(s), or the failure to output new content, only reiterating the prompt. Refusals often contain apologies and explanations of the harms of stereotyping. Tables 6 and 7 show refusals for these prompts by group and by prompt. In general, there is some variance, but it is less correlated with groups, and more on prompt type. Note that GPT does not exhibit refusals with the SCM prompts.
For text generation, we define refusals by outputs that contain only general statements, rather than a specific character or story, or warnings or apologies about not being able to comply with the request. For LLaMA, refusals also include failure to generate any content. Tables 8 and 9 show refusals for text generation, by group and by prompt. The clearest trend is for prompt types. For all LLMs, refusals are seen for Prompts 4 and 5 (Table 4), which request descriptions of an individual. The refusals tended to be apologies, stating the inability to generalize based on gender or sexual orientation (for Prompts 4 and 5). We suspect that these prompts conflict with guardrails intended to avoid discriminatory behavior.
Figure 6: Refusals by group for SCM prompts.
Figure 7: Refusals by SCM prompt type.
Figure 8: Refusals by group for text generation.
Figure 9: Refusals by prompt for text generation. Refer to Table 4 for prompts, which appear in order.